### v2

# Compliance-to-Code: Enhancing Financial Compliance Checking via Code Generation

Abstract
Nowadays, regulatory compliance has become a cornerstone of corporate governance, ensuring adherence to systematic legal frameworks. At its core, financial regulations often comprise highly intricate provisions, layered logical structures, and numerous exceptions, which inevitably result in labor-intensive or comprehension challenges. To mitigate this, recent Regulatory Technology (RegTech) and Large Language Models (LLMs) have gained significant attention in automating the conversion of regulatory text into executable compliance logic. However, their performance remains suboptimal particularly when applied to Chinese-language financial regulations, due to three key limitations: (1) incomplete domain-specific knowledge representation, (2) insufficient hierarchical reasoning capabilities, and (3) failure to maintain temporal and logical coherence. One promising solution is to develop a domain specific and code-oriented dataset for model training. Existing datasets such as LexGLUE, LegalBench, and CODE-ACCORD are often English-focused, domain-mismatched, or lack fine-grained granularity for compliance code generation. To fill these gaps, we present Compliance-to-Code  [^1] https://github.com/AlexJJJChen/Compliance-to-Code, the first large-scale Chinese dataset dedicated to financial regulatory compliance. Covering 1,159 annotated clauses from 361 regulations across ten categories, each clause is modularly structured with four logical elements-subject, condition, constraint, and contextual information-along with regulation relations. We provide deterministic Python code mappings, detailed code reasoning, and code explanations to facilitate automated auditing. To demonstrate utility, we present FinCheck: a pipeline for regulation structuring, code generation, and report generation. Compliance-to-Code establishes a new benchmark for LLM-based compliance automation. Experimental evaluation shows that GLM-4-9B-0414 achieves the best performance on key task in regulation structuring and Qwen3-8B performs the best in compliance code generation tasks.
footnotetext: *These authors contributed equally to this work.footnotetext: †Corresponding Author
1 Introduction
Corporate compliance with financial regulations is regarded as a critical mechanism through which institutions ensure their operations adhere to complex legal frameworks and industry standards (DePamphilis, 2019). Modern financial regulations are notoriously elaborate, featuring multilayered conditions, exceptions, quantitative thresholds, and numerous cross-references, all of which require accurate interpretation and operationalization (Butler and O’Brien, 2019). Traditionally, compliance checking relies heavily on legal, risk, and audit teams, incurring high costs and lengthy processing times. As the volume and complexity of regulations continue to grow, these manual approaches have become increasingly unsustainable (Weirich et al., 2025; Martin et al., 2024).
Regulatory Technology (RegTech) aims to automate compliance management by leveraging advanced computational techniques (Charoenwong et al., 2024). With the emergence of large language models (LLMs), substantial progress has been made in various legal text processing tasks, such as information retrieval, classification, and summarization (Guha et al., 2023; Kant et al., 2025). However, directly employing LLMs for compliance checking poses significant challenges. Due to the inherently black-box nature of LLMs and the non-determinism of their outputs, the generated answers may lack consistency and transparency. By contrast, executable code offers clear, verifiable compliance decisions. This raises a fundamental question: Can LLMs bridge the gap between natural language regulations and executable compliance code, thereby enabling fully automated compliance auditing? To explore this, we conducted preliminary experiments with state-of-the-art LLMs, including Gemini 2.5 Pro (Kavukcuoglu, 2025) and Qwen3-235B (Qwen Team, 2025). As illustrated in Figure 1, our results indicate that existing LLMs still exhibit significant limitations in regulation-to-code tasks:
  • Incomplete Domain Knowledge: LLMs may conflate domain concepts or overlook essential calculation steps. For instance, Gemini 2.5 Pro detected stock distributions by monitoring overall capital changes, ignoring whether such changes resulted strictly from bonus shares, and miscalculated diluted EPS by skipping post-distribution adjustments, resulting in flawed compliance judgments.
  • Shallow Structural Reasoning: LLMs often fail to capture the precise data granularity or semantics required for financial compliance logic. For example, Qwen3-235B designed rule checks around hypothetical or missing input fields instead of leveraging explicit, available columns (e.g., "bonus share ratio"), leading to incomplete or irrelevant condition detection.
  • Temporal and Logical Inconsistencies: Both models demonstrate limitations in tracking event causality and enforcing regulatory constraints over time. Qwen3-235B defaulted to over-simplified constraint satisfactions in the absence of explicit disclosure data, while Gemini 2.5 Pro wrongly applied event flags across all future periods and relied on ambiguous keyword searches in announcements, failing to uphold event-specific, rule-based compliance.
Figure 1: Illustration of error in LLM-generated compliance logic. Code in red color means error.
We attribute these weaknesses primarily to the insufficient domain-specific knowledge of LLMs, and inability to modularize regulatory content, and the absence of appropriate code-oriented training data. Previous research has attempted to address these issues by constructing datasets for improved LLM evaluation in compliance code generation. Existing benchmarks-such as LexGLUE (Chalkidis et al., 2022), LegalBench (Guha et al., 2023), and MAUD (Wang et al., 2023)-focus on tasks like legal text classification, clause extraction, or high-level reasoning. However, they do not modularize financial regulations, and therefore are inadequate for the end-to-end compliance code generation scenario (Küçük and Can, 2025). Other datasets, such as CODE-ACCORD (Hettiarachchi et al., 2025), attempt to formalize condition-action relationships within regulations, but are largely domain-agnostic and not tailored to financial compliance, which frequently relies on specialized terminology. Furthermore, CODE-ACCORD (Hettiarachchi et al., 2025) and similar datasets do not effectively reduce model hallucinations in this domain and are predominantly in English, limiting their applicability to diverse regulatory environments. Consequently, there is a lack of a high-quality, Chinese-language dataset for financial compliance checking-a critical resource for advancing LLMs’ understanding, reasoning, and automation capabilities in regulatory contexts.
To address the lack of automated tools for Chinese financial regulatory compliance, we present the Compliance-to-Code dataset-a large-scale, richly annotated resource tailored for structured understanding and code-based reasoning in this domain.
Dataset Construction and Features: The Compliance-to-Code dataset is built from a systematic analysis of 361 authoritative Chinese financial regulations, spanning ten regulatory categories. We extract and decompose 1,159 representative legal clauses into four fundamental logical components-subject, condition, constraint, and contextual information-to create fine-grained Compliance Units (CUs). We also annotate relationships among CUs to support advanced compositional and cross-reference reasoning. Crucially, we selected CUs amenable to computational assessment, annotated them with executable Python modules, and used these as templates to create 307 coding tasks. Each code example includes detailed Chain-of-Thought reasoning and explanatory comments, enhancing transparency and reproducibility.
FinCheck Pipeline: Automated Compliance Checking. To demonstrate practical utility, we develop FinCheck, a modular end-to-end pipeline leveraging the Compliance-to-Code dataset (see Fig. 3). FinCheck includes: (1) a structure predictor that parses free-text regulations into CUs; (2) a code generator translating CU structures into executable Python code; (3) an information retriever fetching relevant company data; and (4) a report generator producing user-friendly compliance assessments. This pipeline provides traceable, automated compliance checking.
Experimental Results: We benchmark recent LLMs for regulation structuring and code generation tasks using our dataset’s detailed annotations. Qwen3-8B (Yang et al., 2025) achieves the best performance on code generation. With Supervised Fine-Tuning (SFT), GLM-4-9B-0414 notably improves its regulation structuring capabilities.
Summary of Contributions:
1. Compliance-to-Code dataset: We developed the first large-scale Chinese financial regulation dataset structured for code-based compliance reasoning, with 1,159 annotated clauses, 307 Python modules, and reasoning steps.
2. FinCheck pipeline: FinCheck turns the dataset’s structure into a practical, auditable automation solution, supporting robust benchmarking.
3. Comprehensive evaluation: Our experiments provide strong baselines and highlight both challenges and opportunities for code-driven financial compliance in China.
2 Related Works
2.1 RegTech Methods for Financial Compliance
Regulatory Technology (RegTech) leverages computational tools to automate financial compliance (Butler and O’Brien, 2019). Early RegTech solutions relied on deterministic, rule-based systems and expert models (e.g., GEM (Hassan and Logrippo, 2009)) using formal logic and standards like XBRL (Jeyasingh, 2023). While precise and explainable, these symbolic systems struggled with scalability and adaptability amid evolving regulations.
Recent advances in AI, notably Large Language Models (LLMs), have enabled data-driven approaches for parsing regulations and automating compliance workflows (Sun et al., 2025a; Jiang et al., 2025). These neural models offer broader generalization but face challenges with domain specificity, rapid regulatory change, and the complexity of legal logic. Key limitations remain in transparency, interpretability, and auditability-essential for compliance contexts (Wang et al., 2024a).
To address this, hybrid pipelines combining neural models with symbolic reasoning or human oversight have become prominent (Wang et al., 2024a; Sun et al., 2025b). Our work follows this trajectory, presenting a code-centric approach for LLM-driven compliance automation that enhances rigor, reliability, and legal traceability beyond brittle rule-based pipelines.
2.2 Related Datasets and Benchmarks
Despite active research in RegTech, there is a critical lack of datasets mapping real regulatory text to executable compliance logic in the financial sector. Most existing resources address subtasks such as document classification (e.g., LEDGAR (Tuggener et al., 2020), LexGLUE (Chalkidis et al., 2022)), judgment prediction, or QA (e.g., LegalBench (Guha et al., 2023), BillSum (Kornilova and Eidelman, 2019)), but lack annotations for code-level reasoning or logic synthesis.
Datasets closer to financial regulation, like ObliQA (Gokhan et al., 2024) and BSARD (Lotfi et al., 2025), only support question answering or retrieval. CODE-ACCORD (Hettiarachchi et al., 2025) stands out as a logic-annotated resource, but it is limited to building regulations and has narrow domain coverage. Crucially, no existing benchmark, especially in Chinese language systematically couples financial statutes with executable, expert-annotated compliance logic, impeding progress towards true compliance to code. Our work directly addresses this foundational gap.
3 Compliance-to-Code Dataset
3.1 Data Collection and Pre-processing
Compliance-to-Code Dataset is systematically curated from official regulatory documents published by the Beijing Stock Exchange (BSE) [^2] https://www.bse.cn/rule/secnotice_list.html, with a focus on compliance obligations, internal controls, and risk points for publicly listed companies in China. In comparison to the regulations issued by the China Securities Regulatory Commission (CSRC) [^3] http://www.csrc.gov.cn, the BSE provides more detailed and supplementary rules, increasing both the granularity and complexity of regulatory interpretation. We specifically gathered authoritative documents classified as Departmental Rules and Normative Documents, which constitute the regulatory backbone of corporate compliance obligations and reporting requirements. The regulatory corpus reflects all on-force documents as of January 1, 2025, providing a stable but comprehensive snapshot for benchmarking compliance automation.
We identified ten core thematic areas central to corporate compliance program design (see Table 4). Thematic selection was motivated by three criteria: (i) relevance to regulatory risk exposures, (ii) frequency and significance of compliance controls, and (iii) interpretative complexity, thus aligning with both enterprise compliance management frameworks and prevailing regulatory guidance. In addition, considering the dynamic nature of regulatory changes, we prioritized rules with broad applicability and removed highly exceptional or idiosyncratic provisions. For extremely ambiguous regulations-where even expert legal opinion is unavailable-these were deliberately excluded to enhance dataset reliability.
The initial preprocessing stage involved robust legal text extraction, normalization, and standardization, ensuring accurate and complete conversion from diverse official document formats into a unified legal text corpus.
3.2 Data Annotation
Drawing inspiration from RegGPT (Wang et al., 2024b), we observe that regulations are typically composed of six core elements: meta data (entity/property), constraint, condition, measure, scope, and external references. To facilitate more effective translation of regulatory rules into executable code by LLMs, we propose several refinements to this schema. Specifically, we merge the elements of measure and constraint under the unified category of Constraint, relabel external references as Contextual Information, and rename meta data to Subject.
Based on these modifications, we introduce the concept of a Compliance Unit (CU)-the minimal actionable regulatory statement, which can be a rule, requirement, or prohibition. Each CU is structured by four key components: Subject, Condition, Constraint, and Contextual Information (e.g., source, applicability, or exceptions).
To capture compliance logic, we annotate four directed Inter-Unit Relations: refer to (CU references another), exclude (CU negates another), only include (logic applies only if another CU is met), and should include (compliance depends on satisfying another CU). See Appendix F for definitions and examples.
Annotation leveraged a multidisciplinary team in securities law, financial compliance, and audit automation, operating in four stages:
  1. 1.
    CU Identification: Legal/compliance experts segmented regulatory text into CUs, tagging each field and clarifying compliance intent, resolving ambiguities with compliance officers and legal counsel.
  2. 2.
    Inter-Unit Relation Annotation: Experts identified logical relations between CUs.
  3. 3.
    Compliance Logic Archetypes: Domain specialists abstracted 68 distinct logic patterns from annotated CUs and formulated Python-based automation tasks mirroring realistic compliance scenarios. To model regulatory volatility, we varied numerical values and constraints in these code templates, simulating clause amendments and threshold changes.
  4. 4.
    Reasoning and Validation: Each task includes a detailed reasoning trace mapping regulation to code, with dual expert review to ensure accuracy.
We reformulated each programming problem as a code completion task and categorized them into three difficulty levels: simple, medium, and difficult. For Simple problems, the LLM needed to fill in one or two missing lines of code; for Medium problems, three lines; and for Difficult problems, the LLM was tasked with generating the entire solution from scratch. DeepSeek-R1 (Guo et al., 2025) provided candidate extractions and code suggestions, but all outputs underwent independent expert review and adjudication.
Rigorous multi-layered quality assurance encompassed:   (i) Adherence to expert-developed annotation guidelines, (ii) Inter-annotator agreement ($\kappa{=}0.866$ on 15% subset, 180 CUs), (iii) Automated artifact checks and extensive unit testing for code, and (iv) Expert audit of 68 randomly selected automation tasks against source obligations and industry best practice. More details of data annotation are shown in the supplementary material.
3.3 Dataset Statistics
Figure 2: Distribution of Dataset Difficulty Levels, Average Number of Compliance Unit in One Clause and Inter-Unit Relation Types.
The Compliance-to-Code dataset consists of: (1) clean regulatory texts; (2) structured CUs and their inter-unit relations; (3) Python-coded compliance automation tasks; (4) expert-validated COT code generation analyses. The base covers 361 regulations (10 domains) with 53,340 words.
Final annotation produced 1,159 CUs (from 1,226 candidates), and 864 directed inter-unit relations - refer to (61.1%), exclude (2.1%), only include (19.7%), should include (17.1%). Each CU includes averages 0.72 relations, 0.95 subjects, 0.80 conditions, and 0.92 constraints and 0.18 contextual informations. (in Figure 2)
These were mapped to 307 distinct executable compliance automation tasks (Python), each code module has COT reasoning step (average 8 steps/2145 tokens). Tasks stratify by scenario complexity: Simple (70.03%), Medium (11.07%), and Difficult (18.89%).
4 FinCheck: A Structured Compliance Checking Pipeline
FinCheck is a systematic pipeline for automating compliance checks against financial regulations written in natural language. Its core idea is to convert regulatory text into machine-executable code, allowing transparent and reproducible compliance checking on real-world financial data. As shown in Figure 3, FinCheck comprises four modular components:
Structure Predictor: Given a regulatory document in natural language, a fine-tuned large language model (LLM) is employed to extract Compliance Units (CUs) and relations required for compliance assessment. These CUs, along with logical relations among them, are organized into a structured, machine-readable format that makes the regulatory logic explicit.
Code Generator: The structured representation of the regulation (i.e., CUs and their relations) serves as input to a Structured Synthesis LLM (SS-LLM), also fine-tuned for this task, which then generates Python code implementing the compliance checking logic. To enhance reliability and reasoning ability, the LLM is fine-tuned via supervised learning with step-by-step reasoning. The resulting code realizes the regulatory requirements as formal programmatic checks.
Information Retriever: For any given user query, this component retrieves relevant company or client data (based on Wind Information [^4] https://www.wind.com.cn/mobile/EDB/en.html) from structured sources, preparing it as input for the generated compliance-checking code. We utilize Python-based dataframe operations for efficient and flexible data retrieval aligned with the code’s requirements.
Report Generator: The outcome of code execution is summarized using a dedicated LLM, which is prompted to ensure clear, interpretable, and standardized reporting. This report provides end-users with both the compliance result and a transparent trace of the logical reasoning behind the decision.
Figure 3: FinCheck Compliance Checking Pipeline. The framework processes natural language regulations by first using a Structure Predictor to extract key compliance units. These units are then fed into a Code Generator to create verification code. For a specific case, user input triggers an Information Retriever to fetch relevant company data. The generated code is executed with this data, and a Report Generator summarizes the outcome before the final verification result is shown to user.
5 Experimental Setup
5.1 Core Tasks
We investigate three key tasks reflecting the main stages of the compliance checking pipeline:
  • Task 1: Regulation-to-Structure (R2S): Automatically parses raw financial regulatory text into structured Compliance Units (CUs) and their logical relations.
  • Task 2: Structure-to-Code (S2C): Converts (gold or predicted) CUs and relations into functionally correct, auditable Python code implementing compliance logic.
  • Task 3: Regulation-to-Code (R2C): Directly generates executable compliance code from regulatory text, representing the end-to-end challenge.
5.2 Baselines
Since there are currently no models specifically designed to convert regulations into code, we are unable to employ any models directly aimed at Task 2. Furthermore, although RegGPT (Wang et al., 2024b) demonstrates promising performance for Task 1, it is not open-sourced, which limits our ability to use it in our experiments. As a result, we focus our experiments on general LLMs.
Baseline 1: Direct Prompting LLM (DP-LLM). Instruction-tuned LLMs applied to Task 2 and 3; evaluated with zero-shot and one-shot prompts. The prompt is shown in Appendix H.
Baseline 2: Chain-of-Thought LLM (CoT-LLM). As above, but prompts elicit step-by-step reasoning prior to code generation, using demonstration of chain-of-thought examples.
5.3 Dataset Splits
Our dataset comprises 1,159 compliance units (CUs) and 307 compliance automation tasks. We divide the data into training (70%), validation (15%), and test (15%) sets, ensuring stratification by thematic area and task complexity (Simple, Medium, and Difficult). For model training, we exclusively utilize tasks labeled as Difficult, while evaluation is conducted on all three complexity levels within the test set.
5.4 Training Protocols
Base Models: For Task 1, we utilize Qwen3-8B (Qwen Team, 2025), DeepSeek-R1-Distill-Qwen-7B (Guo et al., 2025), GLM-4-9B-0414 (THUDM, 2025), GPT-4.1 (OpenAI, 2025), Gemini-pro 2.5 (Kavukcuoglu, 2025), and Claude-3.7 (Anthropic, 2025). For Tasks 2 and 3, which focus on code generation, we additionally include Qwen2.5-Coder-7B-Instruct (Hui et al., 2024) and DeepSeek-Coder-6.7B-Instruct (Guo et al., 2024) in our evaluation.
Supervised Fine-Tuning (SFT): We train using causal language modeling (next-token prediction), AdamW optimizer, cosine learning rate schedule, batch size 1, weight decay 0.01. The learning rate is set at $1\times 10^{-4}$, epochs $1$, top_p $0.75$, and temperature 00. For LoRA, we use rank 8, $\alpha=32$, and dropout $0.05$. We employ Swift  [^5] https://github.com/modelscope/ms-swift for efficiency. The prompt is shown in Appendix H.
All training and inference are conducted on 16x NVIDIA A800 80GB GPUs for 14 days.
5.5 Evaluation Metrics
For Task 1 (Regulation-to-Structure), evaluation includes CU Field Extraction, assessed using entity-level Precision, Recall, and F1-score for Subject, Condition, and Constraint spans. The temperature was set to 0 for decoding. CU Boundary Identification is evaluated with Precision, Recall, and F1-score for correctly identifying the start and end tokens of each CU. Relation Extraction performance is measured by Precision, Recall, and F1-score for identifying relation types and their CU arguments.
For Tasks 2 & 3 (Code Generation), we primarily evaluate Functional Correctness, measured by Pass@1 and Pass@5. Specifically, Pass@k represents the percentage of generated code samples that successfully compile and pass all associated unit tests within the first k attempts. The temperature was set to 0.6 for decoding. As a secondary metric, we report Code Similarity, quantified by CodeBLEU (Ren et al., 2020), which comprehensively assesses the lexical, syntactic, and semantic similarity between generated code and the reference solutions.
Qualitative and Auditability-focused metrics, primarily used for SS-LLM, include Commentary Traceability, evaluated as the percentage of input CUs and relations explicitly referenced in the generated code comments (manual evaluation on a subset). Explanation Quality is assessed via domain expert ratings (1-5 Likert scale) for clarity, faithfulness, and completeness of generated CoT explanations (for CoT-LLM) and code commentary (for SS-LLM). Finally, Error Analysis involves manual classification of failure modes for incorrect generations on the test set.
6 Results and Analysis
6.1 Task 1 Performance: Advancing Automated Structuring of Complex Regulatory Text
The efficacy of our end-to-end pipeline hinges on accurately transforming convoluted regulatory text into structured, actionable CUs and their inter-unit relations. Table 1 presents a comprehensive evaluation of models fine-tuned on Compliance-to-Code against robust few-shot baselines for Task 1. Our central hypothesis posits that domain-specific fine-tuning, particularly with models adept at capturing nuanced semantics in Chinese legal text is pivotal for this task.
Table 1: Task 1 Performance (Precision, Recall, F1 %) on the Test Set. CU Fields denotes average scores for Subject, Condition, Constraint. CU Identification concerns overall span detection. Best fine-tuned model scores are in bold. SFT indicates Supervised Fine-Tuning.
Model CU Field Extraction CU Boundary Identification
P (%) R (%) F1 (%) P (%) R (%) F1 (%)
Few-shot Performance of Base Models (Task 1: Regulation to Structure, one-shot)
Qwen3-8B $60.74$ $56.88$ $58.74$ $28.57$ $28.57$ $28.57$
GLM-4-9B-0414 $48.01$ $44.96$ $46.44$ $23.21$ $23.21$ $23.21$
GPT-4.1 $64.17$ $63.20$ $63.69$ $40.18$ $40.18$ $40.18$
DeepSeek-V3 71.05 68.01 69.50 49.11 49.11 49.11
Claude-3.7 $65.28$ $64.49$ $64.89$ $41.07$ $41.07$ $41.07$
Fine-tuned Models (Task 1: Regulation to Structure)
GLM-4-9B-0414 (SFT) 72.10 71.27 71.68 37.50 37.50 37.50
Qwen3-8B (SFT) $63.45$ $63.03$ $63.24$ $31.25$ $31.25$ $31.25$
DeepSeek-R1-Distill-Qwen-7B (SFT) $54.97$ $53.28$ $54.11$ $23.21$ $23.21$ $23.21$
Our evaluation demonstrates that domain-adapted supervised fine-tuning is indispensable for structuring complex regulatory text into actionable CUs and inter-unit relations. The fine-tuned GLM-4-9B-0414 achieves 71.7% F1 in CU field extraction, outperforming strong few-shot baselines such as GPT-4.1 by nearly 8 absolute F1 points.
Error analysis localizes major bottlenecks in nuanced relation disambiguation (e.g., should include vs refer to), and multi-hop dependencies. Despite these, the structuring phase delivers a reliable foundation for downstream code synthesis, preserving the majority of regulatory semantics and enabling scalable automation for dense financial texts.
Table 2: Performance of Various Foundation Models before and after SFT for Task 2 (Structure-to-Code, Gold Struct). All models are fine-tuned using our SS-LLM-SFT approach on the Compliance-to-Code dataset. Best results in each category are highlighted in bold.
Model Overall (N=307) Simple (N=215) Medium (N=34) Difficult (N=58)
Pass@1 Pass@5 CodeBLEU Pass@1 Pass@5 CodeBLEU Pass@1 Pass@5 CodeBLEU Pass@1 Pass@5 CodeBLEU
Fine-tuned Model
DeepSeek-R1-Distill-Qwen-7B (SFT) $27.52$ $56.09$ $61.66$ $36.00$ $68.81$ $67.68$ $9.12$ $27.77$ $46.96$ $6.36$ $24.80$ $47.66$
GLM-4-9B-0414 (SFT) $39.72$ $65.94$ $70.68$ $51.11$ $77.12$ $78.60$ $14.71$ $45.12$ $57.70$ $11.45$ $35.92$ $48.34$
Qwen2.5-Coder-7B-Instruct (SFT) $35.80$ $60.15$ 71.57 $48.06$ $76.02$ 79.51 $13.24$ $40.17$ 60.27 $2.73$ $11.62$ $48.09$
Qwen3-8B (SFT) 54.58 80.19 $69.30$ 65.05 88.15 $76.71$ 38.24 68.37 $55.23$ 24.55 56.96 49.61
deepseek-coder-6.7b-instruct (SFT) $35.67$ $64.87$ $68.74$ $46.26$ $78.08$ $76.82$ $11.18$ $32.77$ $55.51$ $10.18$ $34.00$ $45.94$
Base Model
DeepSeek-R1-Distill-Qwen-7B $18.40$ $48.37$ $57.49$ $23.91$ $60.88$ $63.99$ $7.06$ $22.48$ $41.01$ $4.27$ $16.36$ $42.73$
GLM-4-9B-0414 $41.75$ $67.95$ $70.74$ $52.68$ $78.71$ $78.20$ $22.65$ $54.52$ $59.01$ $11.64$ $35.01$ 49.38
Qwen2.5-Coder-7B-Instruct $41.67$ $63.43$ 71.96 $54.83$ $76.99$ 79.91 $17.06$ $43.78$ 60.01 $6.36$ $23.55$ $48.88$
Qwen3-8B 60.03 82.93 $69.81$ 68.10 88.57 $77.52$ 51.18 75.25 $56.24$ 34.55 66.06 $48.63$
deepseek-coder-6.7b-instruct $28.47$ $58.29$ $62.75$ $37.63$ $72.24$ $72.62$ $8.53$ $31.87$ $46.80$ $5.64$ $21.13$ $34.74$
Table 3: Performance Evaluation of Different Models and Approaches on Code Generation Tasks (Model: Qwen3-8B). Gold Struct implies human-annotated structures. Best results in each main section are in bold.
Model / Approach Overall (N=307) Simple (N=215) Medium (N=34) Difficult (N=58)
Pass@1 Pass@5 CodeBLEU Pass@1 Pass@5 CodeBLEU Pass@1 Pass@5 CodeBLEU Pass@1 Pass@5 CodeBLEU
Direct Prompting Baselines (Task 3: Regulation-to-Code, Model: Qwen3-8B)
CoT-LLM (one-shot) 52.29 70.93 69.49 66.42 87.59 77.53 46.76 75.72 56.86 0.56 2.78 46.03
CoT-LLM (0-shot) 52.69 68.34 68.68 67.27 84.44 76.89 45.88 77.02 60.11 0.00 0.00 42.03
DP-LLM (one-shot) 52.86 70.11 69.34 68.03 87.00 77.32 41.47 72.54 56.64 0.74 2.62 46.16
DP-LLM (0-shot) 51.99 70.62 68.60 66.33 88.16 76.93 45.59 73.87 60.12 0.00 0.00 41.40
Direct Prompting Baselines (Task 2: Structure-to-Code, Model: Qwen3-8B, Gold Struct)
CoT-LLM (one-shot) 61.00 82.97 69.76 70.21 89.15 77.22 47.35 75.98 57.53 34.09 63.60 48.68
CoT-LLM (0-shot) 52.08 70.53 69.02 66.71 87.75 77.40 45.59 77.80 60.97 0.00 0.00 41.83
DP-LLM (one-shot) 60.58 82.84 69.83 69.36 88.98 77.63 47.35 73.52 56.88 35.09 65.05 47.90
DP-LLM (0-shot) 53.50 69.69 68.81 68.58 87.35 77.50 46.47 72.85 59.48 0.00 0.00 41.21
Direct Prompting Baselines (Task 2: Structure-Reasoning-to-Code, Model: Qwen3-8B, Gold Struct)
CoT-LLM (one-shot) 78.10 91.28 72.41 84.95 94.21 79.23 62.35 84.34 62.09 62.11 84.57 53.31
DP-LLM (one-shot) 78.13 90.28 73.28 85.40 94.11 79.96 64.41 83.53 63.90 59.39 80.15 54.17
6.2 Tasks 2&3: Structured Regulations Are Transformative for Compliance Code Generation
Effect of SFT and Model Choice.
As shown in Table 2, the effectiveness of supervised fine-tuning (SFT) varies significantly across models. For instance, Qwen3-8B achieves strong performance both as a base model and after SFT, with only marginal differences (e.g., Pass@1: 60.03% base vs. 54.58% SFT). In contrast, other models such as DeepSeek-R1-Distill-Qwen-7B and deepseek-coder-6.7b-instruct benefit more noticeably from SFT. This suggests that the pre-training quality and scale of the foundation model play a critical role in code generation task, and that SFT may yield diminishing returns for already highly capable models.
Impact of Structured Inputs and Reasoning.
Table 3 demonstrates that structured input and explicit reasoning steps substantially improve code generation performance. Moving from direct prompting on raw regulations (Task 3, Pass@1 $\sim$52.9%) to providing human-annotated structures (Task 2, Pass@1 $\sim$61.0%) yields a clear performance boost. Incorporating explicit reasoning (Task 2 Structure-Reasoning-to-Code) further elevates Pass@1 to 78.1%. This progression highlights the importance of decomposing complex compliance requirements into structured representations and guiding the model through intermediate reasoning steps.
Prompting Strategies.
Few-shot prompting and chain-of-thought (CoT) approaches generally outperform zero-shot and direct prompting baselines, especially on more challenging examples. For instance, in the difficult subset, CoT-LLM (one-shot) achieves up to 34.09% Pass@1 (Task 2), compared to near-zero for zero-shot methods. This underscores the value of providing exemplars and reasoning scaffolds, particularly for complex or ambiguous compliance scenarios.
Overall, our findings indicate that while SFT can be beneficial in code generation for some models, the most significant gains in compliance code generation arise from leveraging structured problem decomposition and explicit reasoning in the prompting process. These strategies are especially crucial for handling complex or difficult cases, and their effectiveness is amplified when combined with strong foundation models.
6.3 Auditability, Maintainability, and Granular Error Analysis
Beyond quantitative metrics, the pipeline establishes a new benchmark in auditability and maintainability. Over 92% of generated code artifacts contain directly traceable commentary, supporting regulatory review and compliance traceability beyond black-box LLM generations. Expert evaluation rates the explanatory alignment of our SS-LLM-SFT outputs substantially higher than prompt-engineered baselines (4.5/5.0 vs. 4.1/5.0).
Modularity in generation enables rapid adaptation to regulatory amendments via localized CU updates, a substantial advantage over sequence-level re-prompting approaches. Error decomposition reveals that 45% of failures derive from upstream structure prediction-especially relation errors-highlighting targeted priorities for future research in compositional reasoning and semantic modeling. The remaining errors arise from complex logic mappings, entity normalization, and generic code errors, underscoring the hard boundary between language-level and problem-specific model limitations. For more details of error analysis, please refer to Appendix G.
7 Conclusion
We have introduced the Compliance-to-Code Dataset and a novel structured synthesis pipeline, embodied in our FinCheck system, for executable compliance checking in the complex domain of financial regulation. Our approach, which formalizes regulatory logic as annotated Compliance Units and translates these into verifiable Python code, demonstrates significant gains in functional accuracy (achieving up to 78.1% Pass@1 with gold structures and reasoning steps, interpretability through traceable commentary, and inherent auditability compared to end-to-end LLM prompting. This research not only provides a valuable new resource and strong benchmark results but also opens new avenues for developing safer, more scalable, and reliable RegTech solutions.
Limitations and Future Work: While our approach shows strong results on Chinese financial regulations, extending to other domains and languages may require additional adaptation. Expert annotation remains a bottleneck, though future work could leverage semi-automated or LLM-assisted labeling to improve scalability. Accurately modeling and reasoning about complex legal relationships across documents is still challenging; more advanced graph-based methods or integration of external knowledge bases may help address this. Furthermore, though our current system focuses on translating regulations into compliance logic and granular checks, real-world compliance also involves risk assessment and evaluation of internal controls. Expanding our framework to support these aspects is a promising direction for future research.
Societal Impact: Our approach can streamline compliance and enhance transparency, but should not replace expert judgment in critical situations. We have carefully reviewed and redacted data for release, and recommend that our tools be used with responsible human oversight. For more discussion, please refer to Ethics Statements in Appendix D.
Appendix A License
The Compliance-to-Code dataset will be publicly available and use the CC-BY-NC-4.0 license.
The regulatory documents published by the Beijing Stock Exchange (BSE), used to create the dataset, are publicly available.
Appendix B Accessibility
1. Links to access the dataset and its metadata. (https://huggingface.co/datasets/GPS-Lab/Compliance-to-Code). Link to access code (https://github.com/AlexJJJChen/Compliance-to-Code).
2. The data is saved in both json and csv format, where an example is shown in the README.md file.
3. Logos AI Lab research group will maintain this dataset on the official Github account.
Appendix C Data Usage
The authors bear all responsibility in case of violation of rights.
Appendix D Ethics Statements
The research conducted in this paper adheres in every respect to the NeurIPS Code of Ethics.
We have considered the broader societal impacts of this work. The development of automated tools for financial compliance checking, such as the proposed FinCheck system and the Compliance-to-Code dataset, has the potential for positive societal impact by streamlining compliance processes, increasing transparency, and potentially reducing the costs associated with regulatory adherence. However, we also recognize that such tools should not entirely replace expert human judgment in critical financial and legal situations. Potential negative impacts could arise from over-reliance on automated systems or if the system produces incorrect interpretations, which could lead to erroneous compliance decisions. We recommend that our tools be used with responsible human oversight to mitigate such risks.
In terms of safeguards for the responsible release of data and models, the dataset derived from public regulatory documents has been carefully reviewed. While the primary risk of misuse for a system designed for regulatory compliance is lower than for some other types of AI models (e.g., generative models for media), we emphasize that the generated code and compliance checks should be validated by domain experts before being relied upon for actual compliance decisions. The data and code will be released with clear documentation regarding its intended use and limitations.
All existing assets (e.g., pre-trained models or libraries) used in this paper are properly credited, and their licenses and terms of use have been respected. Any new assets introduced, including the Compliance-to-Code dataset and the FinCheck pipeline code, will be well-documented, and this documentation will be provided alongside the assets to ensure clarity for users and facilitate reproducibility.
This research did not involve crowdsourcing or direct research with human subjects; therefore, IRB approval was not applicable.
The usage of Large Language Models (LLMs) is a core component of the methodology in this research. Specific LLMs employed and their roles in tasks such as regulation structuring and code generation are detailed in the experimental setup section of the paper. This includes information on the base models used and the fine-tuning procedures applied.
Appendix E Acknowledgements
We thank Siyuan Li for leading the dataset construction and experiments, ensuring data quality and experimental accuracy. Jian Chen served as the project lead and contributed to the overall research framework, dataset and experimental design, manuscript preparation, and initial project coordination. Rui Yao contributed to both dataset construction and experimental implementation. Weihua Qiu, Simin Zhang, Chucheng Dong, and Zhiyao Li assisted in dataset creation. Xuming Hu and Peilin Zhou contributed to the revision and optimization of the manuscript. We thank Qipeng Xie and Zixuan Yuan for providing computational resources, with Zixuan Yuan also offering academic guidance. We are grateful to Dr. Yutao Huang (Chinese Academy of Social Sciences) for policy consulting, and to Yuting Qiu for legal advice.
This work is supported by the Guangzhou-HKUST(GZ) Joint Funding Program (No. 2024A03J0630) and NSFC No. 6240074265.
Appendix F Annotation Schema: Compliance Units and Inter-Unit Relations
The foundation of our structured representation is the Compliance Unit, defined as a self-contained, verifiable component of a legal rule. Each Compliance Unit is systematically decomposed into the following constituent fields:
  • Subject: The entity or set of entities legally bound by the rule (e.g., "controlling shareholder," "directors | supervisors | senior management"). We retain precise legal terminology where appropriate and use "|" to denote a disjunction of subjects.
  • Condition: The specific context or trigger scenario under which the constraint applies. This field captures behavioral prerequisites, temporal specifications, quantitative thresholds, and states of third parties. Crucially, exception clauses (e.g., "except for judicial enforcement") are integrated logically within the condition, often transformed into negative constraints.
  • Constraint: The mandatory action, prohibition, or requirement imposed by the rule (e.g., "shall announce 15 days in advance," "must not exceed 3 months," "is prohibited from reducing holdings"). This includes preserving quantitative parameters and explicit modal verbs of obligation ("shall," "must," "is prohibited"). We distinguish between genuine obligations/prohibitions and statements merely conferring rights or permissions; only the former are annotated as constraints unless non-compliance with a permission implies a penalty.
  • Contextual Information: Ancillary details that are necessary for the correct interpretation or execution of the Compliance Unit but do not fit into Subject, Condition, or Constraint. This includes definitions, calculation methodologies, or references to specific annexes or forms. Content that logically belongs in the core fields is excluded.
Regulatory provisions are not usually isolated; they interact logically. We annotate Inter-Unit Relations between Compliance Units to capture these dependencies, enabling the construction of a directed graph representing the legal logic flow. The four primary relation types annotated are:
  • refer to (Reference): A source Compliance Unit requires information from a target Compliance Unit or an external legal document/definition for complete interpretation or execution. This relation type is primarily utilized during the code generation phase to fetch necessary external information.
  • exclude (Exclusion): If the conditions of the source Compliance Unit are met, the target Compliance Unit becomes explicitly inapplicable or is overridden. This relation type modifies the outcome during the post-execution evaluation of the compliance graph.
  • only include (Exclusive Applicability): If the conditions of the source Compliance Unit are met, compliance consideration within the scope of the current regulation is restricted solely to the target Compliance Unit(s), overriding the applicability of other co-located Compliance Units. This relation type also acts during post-execution evaluation.
  • should include (Mandatory Inclusion): The source Compliance Unit explicitly mandates compliance with the constraint(s) of the target Compliance Unit(s). This relation can imply overriding the original condition(s) of the target unit if the source unit’s context directly necessitates the target’s constraint (e.g., a general rule (target) vs. a specific scenario (source) that requires the general rule’s action regardless of the general rule’s specific condition). This relation acts during post-execution evaluation and can influence code logic.
Appendix G Error Analysis
A detailed analysis of failures in the SS-LLM-SFT (Pred Struct) setting reveals critical insights:
  • Upstream Error Propagation (approx. 45% of errors): The most significant error source is inaccuracies from Task 1. Key issues include errors in partitioning regulatory text into its Condition and Constraint components, or failing to fully capture and propagate essential contextual information. Erroneous CU boundary detection can also truncate or conflate conditions.
  • Complex Logic Mapping Failures (approx. 30% of errors): Even with largely correct CUs, mapping highly complex conditional logic, especially involving multiple interacting quantitative thresholds, temporal conditions, or deeply nested exceptions derived from several inter-related CUs, can challenge the code generator. These are typically found in the Difficult problem category.
  • Numerical and Entity Precision (approx. 15% of errors): Financial regulations demand high precision. Errors include slight miscalculations in formulas derived from text, incorrect mapping of specific named entities (e.g., "controlling shareholder" vs. "actual controller" if not perfectly disambiguated in CUs), or off-by-one errors in date calculations.
  • Generic Code Generation Artifacts (approx. 10% of errors): These include minor Python syntax errors, incorrect library usage (though rare with code-tuned LLMs), or slight deviations from the exact expected output format for unit tests, not necessarily reflecting a core misunderstanding of the regulation.
This detailed error breakdown not only highlights current limitations but also provides a clear roadmap for future research. It emphasizes the need for more robust Task 1 models and advanced compositional generalization in the code generation phase. To effectively handle complex multi-CU scenarios, developing stronger relational reasoning capabilities will also be essential.
Appendix H Prompt of Instruction-tuning
Because the length of prompt is long, we only display the structure of the prompt in Figure 4 For more detail, please refer to the dataset.
Figure 4: Prompt of Code Generation.
Table 4: Overview of Compliance Control Domains Covered. The columns # Source clauses, # Compliance Units, and # Document Words indicate the number of original regulatory articles, the total count of extracted structured compliance units, and the total word count of the source documents for each domain, respectively.
Domain/Aspect Compliance Control Focus and Risk # Source clauses # Compliance Units # Document Words
Independent Director Systems Independence requirements, director nomination/integrity review, conflict management. 30 126 5446
Quarterly Reporting Obligations Timeliness, content sufficiency, disclosure controls for periodic reports. 16 32 1888
Equity Incentives and ESOPs Plan approval, risk limits, grantee eligibility, compliance disclosure events. 66 154 8132
Share Repurchase Controls Buyback process, threshold triggers, ban periods, reporting duties. 78 266 11458
Tender Offer Compliance Takeover code triggers, procedural controls, reporting/approval checks. 36 113 4827
Inside Information Management Insider identification, ad hoc event control points, information wall requirements. 25 48 3564
Board Transfer/Listing Change Conditions, procedural controls, risk disclosure for inter-market moves. 21 90 2990
Large Shareholder Transactions Control Sale restriction, blackout periods, reporting triggers, conflict discipline. 29 107 5504
Raised Fund Use Controls Escrow, use restriction, monitoring, board/supervisor review. 27 117 4623
Equity Rights Distribution Dividend/bonus triggers, shareholder fairness, process integrity. 33 106 4908